A car travelling at 80km/h gets a stone caught in one of its tires. Determine an equation that models the height of the stone over time if the radius is 25cm.
1
Expert's answer
2019-05-29T11:17:30-0400

A typical equation for the y-coordinate over time of a point during its circular motion looks like this:


"Y=A\text{sin}(\omega t),"

where "A" - amplitude, "\omega" - angular frequency.

In our case amplitude is the radius of the wheel, the angular frequency can be determined through the velocity and wheel radius as well:


"v=\omega R."

Hence:


"Y(t)=R\cdot\text{sin}\Big(\frac{v}{R}\cdot t\Big),"

make sure you use SI, i.e. you converted km/h to m/s and centimeters to meters:


"Y(t)=0.25\cdot\text{sin}\Big(\frac{80\cdot1000/3600}{0.25}\cdot t\Big),"

so the height over time is

"Y(t)=0.25\cdot\text{sin}(88.89 t)\text{ m}."